OpenStudy (anonymous):

Convergent or divergent? sum_{n = 0}^{infty} frac{ln ^{n} n}{n ^{n} + n}

OpenStudy (anonymous):

oops! question is not clear..

OpenStudy (anonymous):

\[\sum_{n = 1}^{\infty} \frac{\ln ^{n} n}{n ^{n} + n}\] lol the other will be undefined @ n = 0

OpenStudy (anonymous):

i need help!

OpenStudy (anonymous):

i have #1 just need the rest

OpenStudy (anonymous):

please

OpenStudy (anonymous):

this series converges by the ROOT test.. (n th root )

OpenStudy (anonymous):

Solution?

OpenStudy (anonymous):

how do i put it???

OpenStudy (anonymous):

u asked conv. or div.

OpenStudy (anonymous):

yeah but how did you get it well nvm.

OpenStudy (anonymous):